OpenStudy (anonymous):

Solve the equation \frac{ x }{ x+2 } = \frac{ 9 }{ x+9 } Where would i start?

OpenStudy (anonymous):

\[\frac{ x }{ x+2 } = \frac{ 9 }{ x+9 }\]